# Where to find Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

### v2

@@ 3.2 Group-Level Analysis: Later Grokking & Longer Delay for More Difficult Data
4 Memorization-to-Generalization: Analysis of Routing Pathway Dynamics
Grokking is marked by a sudden shift from memorization to generalization long after the training loss has plateaued. Thus, traditional metrics like the training loss provide limited insight into the model’s internal changes leading up to the shift.
In contrast, MoE architectures offer a unique window into the model’s the model dynamics. In an MoE, each input activates only a small subset of experts, racing a discrete sequence of selections (a routing pathway) across layers. This routing pathway effectively externalizes part of the model’s computation, making internal decision-making directly observable.
@@ 4 Memorization-to-Generalization: Analysis of Routing Pathway Dynamics
In this section, we analyze the evolution of pathway patterns during grokking and introduce two metrics to quantify these changes (Figure 3): (i) the similarity of pathways across different inputs, and (ii) the consistency of a single input’s pathway. Importantly, all metrics are computed on the grokking subsets $\smash{\{\mathcal{D}_{\text{grok}}^{(t)}\}}$ identified in Section 3.2, which consist of training samples that have achieved stable, low loss at different checkpoints. Focusing on these grokking samples allows us to isolate and study the internal transition from memorization to generalization, beyond mere loss convergence.
We show that as training progresses, pathways evolve from initially random and instance-specific to increasingly structured and shared across samples, providing a concrete, mechanistic signal of the memorization-to-generalization transition.
@@
Figure 3: Pathway Complexity Metrics to Monitor Grokking. Left: Pathway similarity between samples is measured by edit distance on their sequences of expert choices. Right: Pathway consistency quantifies the smoothness of expert transitions between subsequent layers via cosine similarity of their weighted expert embeddings.
@@ 4.1 Pathway Distance and Knowledge Sharing between Training Samples
One hypothesis is that as the model begins to generalize, it moves from highly individualized, input-specific computation toward shared, structured computation across similar inputs. This shift would be reflected in increasing pathway similarity—the convergence of routing patterns—signaling the emergence of shared mechanisms for solving related tasks. We test this by examining how expert pathway similarity evolves during pretraining. Each input’s pathway is the ordered sequence of selected experts through the $L$ MoE layers. Specifically, for input $x_{i}$, we define its pathway $s_{i}$ as:
$s_{i}=\mathrm{concat}(e_{1}^{(i)},\,e_{2}^{(i)},\,\dots,\,e_{L}^{(i)})$
where $e_{\ell}^{(i)}$ is the ordered list of expert indices at layer $\ell$ for input $x_{i}$, as a comma-separated string (e.g., ‘3,1,5’). These are joined across layers with hyphens (e.g., ‘3,1,5-…-9,1’). Expert selection dynamically includes top experts whose cumulative routing weights exceed a predefined threshold.
We quantify pathway similarity using sequence edit distance:
$D_{\mathrm{path}}(s_{i},s_{j})=\mathrm{EditDistance}\big{(}s_{i},s_{j}),$
@@
where $\mathrm{EditDistance}(\cdot,\cdot)$ is the standard Levenshtein edit distance, counting the minimal number of insertions, deletions, or substitutions needed to transform one sequence into the other. Thus, edit distance captures both local and global differences across the entire routing sequence—accounting for mismatches in expert selection, variations in sequence length, and shifts in expert ordering. We track the average pairwise edit distance across samples over pretraining.
Overall pattern. Our analysis, shown in Figure 4, reveals a clear trajectory in pathway dynamics. Early in pretraining, most inputs follow nearly identical routes through the experts, resulting in low edit distance $D_{\mathrm{path}}$ across samples. As training progresses, however, pathways begin to diverge: each sample “carves out” its own route, increasing $D_{\mathrm{path}}$ between samples. Strikingly, this divergence reverses once convergence sets in. Inputs with similar semantic content are increasingly routed through similar expert sequences, reflecting an emergence of shared pathways. This convergence signals that the model is discovering common internal computations for related inputs—a concrete marker of generalization.
@@
Layerwise pattern. Interestingly, this transition is highly nonuniform across depth, as shown in Figure 5. Shallow layers exhibit the steepest declines in $D_{\mathrm{path}}$ at the grokking point, while deeper layers show smaller reductions. Notably, the final layer reverses this trend, with $D_{\mathrm{path}}$ increasing post-grokking. These results suggest that grokking reorganizes the network in a depth-specific manner: early layers converge toward universal pathways, whereas later layers maintain—or even increase—routing flexibility, balancing shared representations with task-specific specialization.
Figure 4: Pathway edit distance and training loss of two data groups (converged at $125k$ and $245k$ steps) over pretrianing steps across four domains. While training loss plateaued early, the edit distance between pathways keeps declining beyond convergence, indicating a declining complexity of internal memorization and a continuous transition from memorization to generalization.
@@
Figure 5: Layer-wise evolution of pathway edit distance during pretraining. The overall distance drops from early to later layers, indicating higher complexity in early layers. For each layer, the distance initially rises, then sharply declines at grokking onset. The shift is depth-dependent—early layers’ expert choices rapidly converges, later layers change more gradually, and the final layer diversifies after grokking.
@@ 4.2 Pathway Consistency and Complexity for Single Samples
We next investigate how the routing complexity of individual inputs evolves during pretraining. Intuitively, more streamlined and consistent expert selection across layers may signal that the model has discovered a generalizable processing strategy. To capture this, we propose a metric for single-sample pathway consistency, measuring the smoothness of expert transitions across consecutive layers.
In a typical MoE architecture, the router at each layer $\ell$ is implemented as a single-layer MLP—a linear transformation followed by a softmax—that computes routing scores over $K$ experts. The weight matrix of this router is denoted $\mathbf{W}_{\ell}\in\mathbb{R}^{K\times d}$, where $K$ is the number of experts and $d$ is the embedding dimension. We define the expert embedding for expert $k$ at layer $\ell$ as: $\mathbf{v}_{\ell}^{k}:=\mathbf{W}_{\ell}[k],\text{for }k=1,\ldots,K.$ Each $\mathbf{v}_{\ell}^{k}$ captures the router’s parameterized representation of expert $k$, and can be interpreted as the anchor vector used to compute routing scores by comparing it against input features (e.g., via dot product or similarity). Although these embeddings are fixed across inputs, their relative geometry governs expert selection during inference. For a given input $x_{i}$, we define its routing embedding as the weighted sum:
@@
where $g_{\ell}^{(i,k)}$ is the routing weight assigned to expert $k$ for input $x_{i}$ at layer $\ell$. Although experts are independently parameterized across layers, we assume their embeddings $\{\mathbf{v}_{\ell}^{k}\}$ are comparable across layers, enabling structural analysis of routing behavior.
@@
Figure 6 tracks the evolution of pathway consistency and training loss across four tasks. A key finding is that even after the loss stabilizes, pathway consistency continues to improve—revealing ongoing refinement in the model’s internal routing. This suggests that smoother, more coherent cross-layer transitions emerge well beyond the point of loss convergence, positioning pathway consistency as a valuable lens for probing deeper model organization beyond prediction error.
Figure 6: Pathway consistency and training loss of two data groups (converged at $490k$ and $615k$ steps) over pretrianing steps across four domains. While training loss converges early, pathway consistency continues to improve, revealing continuous refinement and reducing complexity of expert choices beyond memorization, which indicates a transition to generalization.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
Table 1: Correlation between pathway metrics and generalization (test accuracy) across four domains, compared with training/validation loss for monitoring generalization. $p$-value: $p<0.1$, $p<0.05$, $p<0.01$, $p<0.001$, $p>0.1$, $p>0.5$, $p>0.8$. Preferred correlation: positive, negative.
@@
Figure 7: Correlation between generalization-monitor metrics and test accuracy, and linear regression on five checkpoints. Pathway edit distance is negatively correlated with test accuracy, while pathway consistency is positively correlated—both aligning with expectations. In contrast, training loss and its moving average show weak or even misleading correlations with test accuracy, underscoring that lower training loss does not reliably indicate better generalization.
Given the strong link between routing dynamics and generalization, we ask: To what extent do our pathway metrics predict test performance across domains? To answer this, we compute both Pearson and Spearman correlations between test accuracy and two metrics across training checkpoints: (1) pairwise pathway similarity, measured by normalized edit distance, and (2) single-sample pathway consistency, measured by cross-layer cosine similarity.
We define the domain-level grokking point as the first checkpoint after which test accuracy for each domain exhibits a sustained upward trend (Figure 1), marking the onset of meaningful generalization. Our correlation analysis focuses on checkpoints after this grokking point—excluding the initial random-accuracy phase—to isolate the mechanisms driving generalization.
Table 1 summarizes the correlation results across four tasks. Pathway consistency (sample-wise) shows exceptionally strong positive correlations with test accuracy—often exceeding $0.97$ and highly significant—indicating that coherent, stable routing across layers is a key marker of generalization. In contrast, pairwise pathway similarity consistently yields strong negative correlations (around -$0.93$), suggesting that as generalization improves, inputs tend to follow increasingly similar expert routes. By comparison, training and validation loss metrics exhibit weaker and less consistent correlations. Notably, while moving-average training loss shows relatively high correlation in certain domains (e.g., Code, Domain-Specific), its direction is inconsistent with expectation: since loss is minimized during training, we would expect a negative correlation with test accuracy—not the positive one observed here. This mismatch further underscores that conventional loss metrics fail to reflect the internal transition from memorization to generalization.
@@ 4.4 Theoretical Connections & Explanations
We establish a fundamental connection between expert routing patterns and model generalization through spectral analysis of the routing kernel geometry. Our theory identifies the effective complexity of the routing mechanism as the key determinant of MoE generalization.
Let $\bm{H}^{*}$ denote the routing kernel Gram matrix with entries $\bm{H}^{*}_{ij}=\Theta_{\text{route}}(x_{i},x_{j})$, where the routing kernel $\Theta_{\text{route}}$ captures parameter sensitivity of expert selection.
@@
Under NTK regime with $K$ experts, with probability $1-\delta$ over $n$ samples:
@@
This explains the grokking phenomenon: When expert activations transition from random to structured, the effective dimension $\text{Tr}(\bm{H}^{*}(\bm{H}^{*}+\lambda I)^{-1})$ collapses, producing sudden drops in generalization error.
For a detailed theoretical setup, assumptions, and proof of Theorem 4.1, please refer to Appendix A.
@@ 5 Conclusion
This paper presents the first empirical evidence of grokking in large-scale LLM pretraining, showing that generalization emerges asynchronously across domains—often well after training loss has converged. Through an analysis of routing dynamics in a 7B-parameter MoE model, we propose two novel pathway-based metrics that accurately track generalization without requiring finetuning or external validation sets. Our findings reveal a transition from memorization to generalization, marked by increased pathway similarity and consistency, and supported by theoretical connections between routing structure and generalization bounds. These insights offer a foundation for more transparent and reliable monitoring of generalization in large-scale models.
Limitations. Our study is limited to a single model architecture due to checkpoint availability. Moreover, the theoretical analysis is restricted to a one-layer MoE under NTK assumptions, which may not extend directly to deeper or more heterogeneous architectures.
Appendix A Generalization Bound for MoE with Routing Kernel
A.1 Model Setup and Assumptions

### v3

@@ 3.2 Asynchronous Memorization, Delayed Generalization, & Local Grokking
4 Monitor Memorization-to-Generalization Transition by Training Dynamics of Routing Pathways
As shown in Section 3, generalization in LLMs emerges well after memorization, suggesting that internal reorganization underlies this transition. Such dynamics are difficult to track in dense architectures, where all neurons are simultaneously involved. MoEs simplify this process: by organizing computation into experts, each input activates only a subset of experts, forming a discrete routing pathway across layers (Figure 3(a)). These pathways reveal how computation is allocated and reorganized during pretraining, providing a mechanistic view of the memorization-to-generalization transition.
@@ 4 Monitor Memorization-to-Generalization Transition by Training Dynamics of Routing Pathways
In this section, we analyze the evolution of routing pathway during pretraining and introduce two metrics to quantify these changes (Figure 3): (i) the similarity of pathways across different inputs (Section 4.1), and (ii) the consistency of a single input’s pathway (Section 4.2). Importantly, all metrics are computed on pretraining samples that have already been memorized, i.e., those whose training objective has stabilized at low values beyond specific checkpoints, and trace how their internal routing continue to evolve from memorization to generalization. We then assess these metrics as robust generalization monitors (Section 4.3) and conclude by establishing a spectral framework connecting routing geometry to generalization bounds (Section 4.4).
@@
Figure 3: Pathway Complexity Metrics to Monitor Grokking. (a) Pathway similarity between samples is measured by edit distance on their sequences of expert choices across layers. (b) Pathway consistency quantifies the smoothness of expert transitions between subsequent layers for the same sample by cosine similarity of their weighted expert embeddings.
@@ 4.1 Pathway Distance and Knowledge Sharing between Training Samples
Building on prior findings that grokking coincides with the emergence of structured internal mechanisms [13, 18, 26], we hypothesize that as the model begins to generalize, it transitions from highly individualized, input-specific computation toward more shared, structured computation across related inputs. This shift should manifest as increasing pathway similarity—the convergence of routing patterns across similar samples—indicating that the model is developing common mechanisms for solving related tasks.
We test this by examining how the similarity of expert pathway evolves during pretraining. Each input’s pathway is the ordered sequence of selected experts through the $L$ MoE layers. Specifically, for input $x_{i}$, we define its pathway $s_{i}$ as: $s_{i}=\mathrm{concat}(e_{1}^{(i)},\,e_{2}^{(i)},\,\dots,\,e_{L}^{(i)})$, where $e_{\ell}^{(i)}$ denotes the ordered list of expert indices at layer $\ell$, obtained by ranking experts according to the mean routing weights across all tokens for input $x_{i}$. To determine which experts are included, we first rank experts based on their routing weight and then select top-ranked experts until their cumulative routing weight exceeds a predefined threshold. The weight threshold selects the experts that actually drive each input’s computation, rather than picking a fixed number of experts. The selected experts are then concatenated as a comma-separated string (e.g., ‘3,1,5’) which are later joined across layers with hyphens (e.g., ‘3,1,5-…-9,1’).
@@
We use the Levenshtein edit distance $D_{\mathrm{path}}(s_{i},s_{j})=\mathrm{EditDistance}\big(s_{i},s_{j})$, which counts the minimum insertions, deletions, or substitutions required to align two pathways, to measure pathway similarity. This metric captures both local deviations in expert choice and global shifts in pathway structure, reflecting differences in selection, length, and ordering. We monitor the average pairwise distance across samples throughout pretraining to track the emergence of shared routing.
Overall pattern. Our analysis, shown in Figure 4, reveals a clear trajectory in pathway dynamics. Early in pretraining, most inputs share nearly identical routes, producing low edit distance $D_{\mathrm{path}}$. As the model memorizes individual inputs, pathways diverge and $D_{\mathrm{path}}$ rises. Crucially, $D_{\mathrm{path}}$ decreases after memorization, which indicates that semantically related inputs then converge into similar routing sequences, signaling the emergence of shared pathways that support generalization.
@@
Layerwise pattern. Figure 5 reveals that the memorization-to-generalization transition is highly depth-dependent. Shallow layers show the steepest decline in $D_{\mathrm{path}}$ after memorization, indicating rapid convergence to shared pathways. Deeper layers exhibit smaller reductions, and the final layer even shows an increase in $D_{\mathrm{path}}$ post-memorization. These patterns suggest a depth-specific reorganization: early layers consolidate universal representations, while later layers retain or enhance routing flexibility, balancing shared computation with task-specific specialization.
Figure 4: Pathway edit distance and pretraining objective of two data groups (memorized at $125k$ and $245k$ steps) during pretraining across four domains. The global panel (leftmost) fits the overall quadratic trend over all domains, while domain panels fit separate trends after the corresponding convergence step. Despite early training objective plateauing, pathway edit distance continues to decline, indicating a declining complexity of internal memorization.
@@
Figure 5: Layer-wise evolution of pathway edit distance during pretraining for the data group memorized at $245k$ step. Overall distance decreases from early to later layers, indicating higher complexity in early layers. After pretraining objective convergence, early-layer routing among similar samples simplifies quickly, later layers adapt more gradually, and the final layer diversifies after memorization.
@@ 4.2 Pathway Consistency and Complexity for Single Samples
We next analyze how the routing complexity of individual inputs evolves during pretraining. Our hypothesis is that increasingly streamlined and consistent expert selection across layers reflects the reuse of shared pathways, signaling the emergence of generalizable processing strategies. To capture this, we define single-sample pathway consistency, which measures the smoothness of expert transitions across consecutive layers.
In a typical MoE, the router at layer $\ell$ is a single-layer MLP—a linear transformation followed by a softmax—producing scores over $K$ experts, with weight matrix $\mathbf{W}_{\ell}\in\mathbb{R}^{K\times d}$. We define the expert embedding for expert $k$ as $\mathbf{v}_{\ell}^{k}:=\mathbf{W}_{\ell}[k],k=1,\dots,K,$ which serves as an anchor vector to compute routing scores. For input $x_{i}$, its routing embedding is the weighted sum:
@@
where $g_{\ell}^{(i,k)}$ is the average routing weight assigned to expert $k$ across all tokens of input $x_{i}$. Although experts are independently parameterized, the strong correlations between adjacent layers in deep networks [6] and MoEs [12] suggest that $\{\mathbf{v}_{\ell}^{k}\}$ can be comparable across neighboring layers, supporting structural analysis of routing behavior.
@@
Figure 6 tracks the evolution of pathway consistency and pretraining objective across four domains. A key finding is that even after the objective stabilizes, pathway consistency continues to improve—revealing ongoing refinement in the internal routing of the model. This suggests that smoother, more coherent cross-layer transitions emerge well after memorization, positioning pathway consistency as a valuable lens for understanding generalization beyond what the training objective alone captures.
Figure 6: Pathway consistency vs. training objective on four domains during pretraining. The global panel (leftmost) shows that the path consistency consistently improves over all data, while each domain panel shows the same trend on two data groups (memorized at $490k$ and $615k$) from each domain after the training objective converges. This reflects a progressively smoother transition of experts between consecutive layers after memorization.
@@ 4.3 Pathway Complexity Metrics as Monitors of Generalization
Table 1: Correlation between pathway complexity metrics and generalization (test accuracy) across four domains, with the comparison to training/validation loss. $p$-value: $p<0.1$, $p<0.05$, $p<0.01$, $p<0.001$, $p>0.1$, $p>0.5$, $p>0.8$. Preferred correlation: positive, negative.
@@
Figure 7: Correlation of the two pathway complexity metrics with benchmark performance. Compared to the training objective and its moving average, both the pathway edit distance and consistency are highly correlated with the test accuracy, though they are also computed on training data.
Given the strong link between routing dynamics and generalization, we ask: To what extent do our pathway metrics predict test performance across domains? To answer this, we compute Pearson and Spearman correlations between downstream performance and two metrics across pretraining checkpoints. Our correlation analysis focuses on checkpoints after domain-level accuracy begins its sustained rise (shown in Figure 1), thereby isolating the mechanisms that drive generalization beyond the random-accuracy phase.
Table 1 summarizes the correlation results across four tasks. Pathway consistency (sample-wise) shows exceptionally strong positive correlations with test accuracy—often exceeding $0.97$ and highly significant—indicating that coherent, stable routing across layers is a key marker of generalization. In contrast, pairwise pathway similarity consistently yields strong negative correlations (around -$0.93$), suggesting that as generalization improves, inputs tend to follow increasingly similar expert routes. By comparison, as shown in Figure 7, training and validation loss metrics exhibit weaker and less consistent correlations. Notably, while moving-average training loss shows relatively high correlation in certain domains (e.g., Code, Domain-Specific), its direction is inconsistent with expectation: since loss is minimized during training, we would expect a negative correlation with test accuracy—not the positive one observed here. This mismatch further underscores that conventional pretraining metrics fail to reflect the internal transition from memorization to generalization.
@@ 4.4 Theoretical Connections & Explanations
We connect expert routing patterns with model generalization via the complexity of routing, formalized through the effective dimension of the routing kernel. This effective dimension will serve as the key determinant of MoE generalization.
Concretely, $\bm{H}^{*}$ is the Gram matrix of the routing kernel with entries $\bm{H}^{*}_{ij}=\Theta_{\text{route}}(x_{i},x_{j})$, capturing similarity between inputs induced by both routing overlap and expert functions. The effective dimension is defined as $\text{Tr}(\bm{H}^{*}(\bm{H}^{*}+\lambda I)^{-1})$, measuring the complexity of the function space induced by routing.
@@
Under NTK regime with $K$ experts, with probability $1-\delta$ over $n$ samples, where $C_{1},C_{2}$ are NTK-dependent constants and $\lambda$ is the regularization parameter:
@@
This provides a lens on grokking: as routing evolves from random to structured, the effective dimension collapses, triggering improvements in generalization. Intuitively, this collapse corresponds to a simplification of routing pathways: the kernel spectrum becomes more concentrated, reducing the hypothesis space and mitigating overfitting. The full theoretical setup, assumptions, and proof are detailed in Appendix C, while Appendix C.4 empirically confirms that effective dimension is tightly aligned with routing edit distance, establishing pathway complexity as a key driver of MoE generalization.
@@ 5 Conclusion
This paper provides the first empirical evidence that, in large-scale LLM pretraining, generalization can emerge well after memorization has been achieved. Through an analysis of routing dynamics in a 7B-parameter MoE model, we propose two novel pathway-based metrics that accurately track generalization without requiring finetuning or external validation sets. Our findings reveal a transition from memorization to generalization, marked by increased pathway similarity and consistency, and supported by theoretical connections between routing structure and generalization bounds. These insights offer a foundation for more transparent and reliable monitoring of generalization in large-scale models. Future work will extend these pathway-based metrics beyond MoE, by constructing analogous “virtual pathways” in dense models, moving toward a unified framework for tracking generalization in foundation models.
Appendix A Dataset Details
We evaluate OLMoE on four domains used during pretraining: math, code, commonsense, and domain-specific (scientific text). For each domain, we construct three sets:
  • Pretraining sample set: 10,000 examples drawn from OLMoE’s pretraining data.
  • Test sample set: 10,000 examples from a benchmark dataset in the same domain, following OLMoE’s categorization where available.
  • Validation set: a disjoint dataset used solely to compute validation loss, enabling correlation analysis between in-domain learning and test generalization.
For the domain-specific case, we filter the scientific pretraining data using the keyword “Organic Chemistry”, aligning the training distribution with the “college_chemistry” benchmark in MMLU.
Table 2 summarizes all dataset sources. All pretraining sets are directly drawn from OLMoE’s pretraining corpora; test benchmarks are standard evaluation datasets; validation sets are only used for correlation analysis.
Table 2: Datasets used in our domain-level evaluation. Pretraining samples are from OLMoE corpora, test benchmarks are used for evaluation, and validation sets are disjoint corpora used only for loss–accuracy correlation. All datasets accessed via HuggingFace.
Domain Pretraining Source Test Dataset Validation Dataset
Math open-web-math/open-web-math hkust-nlp/dart-math-pool-math qwedsacf/competition_math
Code bigcode/starcoderdata evalplus/humanevalplus newfacade/LeetCodeDataset
Commonsense emozilla/dolma-v1_7-books tau/commonsense_qa kowndinya23/flan2021-commonsense
Domain-Specific blester125/mediawiki-dolma(filtered “Organic Chemistry”) cais/mmlu(“college_chemistry” subset) chemNLP/chemistry-bookshelves-merged
Appendix B Experiment Details
Instruction Tuning. The raw OLMoE checkpoints exhibit weak instruction-following ability, making direct evaluation of generalization behavior unreliable—particularly in text-based domains such as commonsense reasoning or code completion. To address this, we apply light instruction tuning using LoRA [7] to make the pretrained model follow prompts in a human-readable way. This ensures we can evaluate generalization while preserving the original pretraining behavior.
Key configurations of LoRA include:
  • Target modules: q_proj, k_proj, v_proj
  • LoRA rank: 32
  • LoRA dropout: 0.1
  • Learning rate: 5e-5
  • Batch size: 4
  • Max sequence length: 2048
Applying Min-K%++ for Training-Test Separation. To maintain a rigorous separation between the pretraining and test datasets, we employ the Min-K%++ membership inference attack [28], a state-of-the-art, reference-free method designed to detect whether specific data samples are part of a model’s training corpus.
Specifically, we apply the Min-K%++ method to the test dataset to identify samples that the model is most likely to have encountered during training. Determining an optimal threshold for classification is challenging due to variability in model behaviors and data distributions. To address this, we adopt a conservative approach by removing the top 10% of test samples with the highest Min-K%++ scores, thereby minimizing the risk of training data contamination in the test set.
Appendix C Generalization Bound for MoE with Routing Kernel
C.1 Model Setup and Assumptions